# Towards Human-like Preference Profiling in Sequential Recommendation

### v1

@@ Abstract
Sequential recommendation systems aspire to profile users by interpreting their interaction histories, echoing how humans make decisions by weighing experience, relative preference strength, and situational relevance. Yet, existing large language model (LLM)-based recommenders often fall short of mimicking the flexible, context-aware decision strategies humans exhibit, neglecting the structured, dynamic, and context-aware mechanisms fundamental to human behaviors. To bridge this gap, we propose RecPO, a preference optimization framework that models structured feedback and contextual delay to emulate human-like prioritization in sequential recommendation RecPO exploits adaptive reward margins based on inferred preference hierarchies and temporal signals, enabling the model to favor immediately relevant items and to distinguish between varying degrees of preference and aversion. Extensive experiments across five real-world datasets demonstrate that RecPO not only yields performance gains over state-of-the-art baselines, but also mirrors key characteristics of human decision-making: favoring timely satisfaction, maintaining coherent preferences, and exercising discernment under shifting contexts. Code: https://anonymous.4open.science/r/RecPO-020A/
Towards Human-like Preference Profiling in Sequential Recommendation
@@ 1 Introduction
In modern social media, recommender systems are ubiquitous in shaping user experiences by delivering personalized content across various online platforms Schafer et al. (1999); Sarwar et al. (2000); Ma et al. (2008); Davidson et al. (2010); Jamali and Ester (2010); Fan et al. (2019); Deldjoo et al. (2020); Ouyang et al. (2024, 2025).
Figure 1: Human-like decision behavior in preference-based choices, involving fine-grained trade-offs among reward, delay, effort, and risk.
Sequential recommendation, as a specialized form, predicts the next item a user is likely to interact with based on his/her historical behaviors. This form plays a crucial role in the real world, such as movie recommendations on streaming platforms like Netflix, or curating the next song on Spotify based on listening history. Sequential recommenders Sun et al. (2019); Tang and Wang (2018); Chang et al. (2021) originally leverage neural models such as RNNs Hidasi (2016) and Transformers Kang and McAuley (2018). These models encode user/item IDs and contextual features (e.g., item descriptions, interaction timestamps) for sequential pattern modeling. More recently, the emergence of large language models (LLMs), has sparked interest in their use for sequential recommendation Harte et al. (2023); Li et al. (2023); Yang et al. (2024); Bao et al. (2023); Zhang et al. (2023), owing to their broad world knowledge and advanced reasoning capabilities.
As central to sequential recommendation, user preference profiling reflects how humans adapt choices based on prior experience, personalized relevance, and evolving goals Astington and Jenkins (1995). However, current LLM-based solutions often fall short in emulating the flexible decision patterns exhibited by humans. Existing preference modeling approaches, such as direct preference optimization (DPO) Rafailov et al. (2024) and its derivatives Chen et al. (2024); Meng et al. (2024); Amini et al. (2024), rely on static pairwise comparisons that abstract away the layered and situational nature of real-world preferences. This modeling overlooks the hierarchical structure of human preference (e.g., mild vs. strong affinity) and fails to account for factors like satisfaction delay, decision effort, and perceived risk, as illustrated in Figure 1. For instance, a moderately preferred item that is immediately relevant may be favored over a highly rated item that becomes relevant much later. Such prioritization, common in human decision-making, is rarely reflected in LLM-based systems, limiting their alignment with real user decision dynamics.
To explore this, we begin with a proof-of-concept experiment to demonstrate that leveraging comprehensive and structured user feedback substantially improves preference modeling in sequential recommendation. Building on this insight, we then introduce RecPO, a framework designed to enhance LLM-based recommenders by emulating human-like decision patterns observed in structured and context-aware feedback. RecPO is distinguished by four key components:
  1. (i)
    It retains all user interactions, including those with negative feedback, to build a fuller picture of preferences and aversions;
  2. (ii)
    It incorporates structured preference inferred from explicit or implicit feedback;
  3. (iii)
    It calibrates preferences data using satisfaction immediacy reflected in humans; and
  4. (iv)
    It generates adaptive reward margins to guide alignment based on both preference hierarchy and temporal context.
Through experiments across five datasets with either implicit or explicit feedback, we show that RecPO more faithfully mirrors human-like behavior: consistently promoting preferred items, suppressing dispreferred ones, and adapting to context that aligns with real-world user decision patterns.
2 Related Work
Sequential recommendation models temporal user preferences in interaction sequences. Early methods adopt structures such as recurrent neural networks (GRU4Rec Hidasi (2016)) and self-attention mechanisms such as SASRec Kang and McAuley (2018). Recent advances integrate LLMs for their rich semantic understanding and contextual reasoning capabilities Liao et al. (2024); Bao et al. (2023); Geng et al. (2022); Yuan et al. (2023).
LLM preference alignment techniques aim to align language models’ outputs with human preferences, have inspired significant advancements beyond general-purpose tasks such as recommendations. Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) and DPO Rafailov et al. (2024) suggest fine-tuning based on human preference data. Building on DPO, methods like IPO Azar et al. (2024), CPO Xu et al. (2024), KTO Ethayarajh et al. (2024), SimPO Meng et al. (2024), and ODPO Amini et al. (2024) further refine alignment with improved model efficiency and robustness. Most recently, S-DPO adapts alignment for user-item sequences, optimizing personalization by comparing with list-wise negative items. We provide a more detailed related work in Appendix D.
Our approach differs from previous efforts in that we calibrate structured implicit or explicit preference feedback based on satisfaction delay to refine preference alignment margins, without obtaining them computationally Amini et al. (2024).
@@ 3 Preliminaries
where $\pi_{\text{ref}}$ is commonly set to $\pi_{\text{SFT}}$, $\pi_{\theta}$ is the aligned model, and $\beta$ is a hyperparameter.
Building upon DPO, a recent effort named S-DPO Chen et al. (2024) has been proposed specifically for LLM-based recommenders. They pair each positive item with multiple negative items generated by random sampling as preference data, and revise the alignment objective as:
@@
Figure 3: Illustrations for DPO and our framework: DPO assigns rigid preference margin across pairwise preference data, while ours adaptively adjusts preference margin ($\delta$) based on preference structures and contextualization.
4 Comprehensive and Structured Feedback Grounds Preference Profiling
We design a proof-of-concept experiment to assess how comprehensive and structured preference feedback, captured by user-item ratings, impacts predicting the next preferred item. We devise four-tier input configurations that progressively integrate preference signals:
  1. (i)
    Filtered Items: Excluding negative feedback items and no explicit ratings are provided, mimicking S-DPO’s setup;
  2. (ii)
    Full Items: Retaining all historical items, yet no explicit ratings are provided;
  3. (iii)
    Filtered Items + Rating: Providing explicit ratings yet excluding negative-feedback items;
  4. (iv)
    Full Items + Rating: Retaining all items and their corresponding explicit ratings.
We fine-tune LLaMA3-8B on MovieLens and Amazon-Books ( described Section 6.1) using the four input configurations. The experimental results are reported using Hit Ratio@1 (see Section 6.1, where higher values indicate better performance) and are shown in Figure 2.
We observe that maintaining comprehensive and structured feedback in the form of ratings consistently improves performance. While it seems counterintuitive to include negatively interacted items when the task is to predict the next favorable items, having structured feedback in the user history proves beneficial, as aversion modeling is crucial for constructing a more accurate user preference profile. Notice that the Full Items (without ratings) underperforms the Filtered Items, as the absence of explicit annotations for negative items introduces noise into the learning process. These results highlight that structured feedback helps resolve ambiguity, enabling LLMs to differentiate between preferences and aversions.
@@ 5 Methodology
In this section, we first lay out the prompt design that establishes the foundation for preference modeling in LLM-based recommendations. We then introduce RecPO, a novel preference optimization framework for sequential recommendation that dynamically calibrates reward margins between pairwise preference data based on contextualized and structured preference feedback, shown in Figure 3.
5.1 Comprehensive and Structured Preference Feedback as the Foundation
Unlike existing approaches Liao et al. (2024); Chen et al. (2024) which remove items with negative feedback from interaction histories to construct homogeneous sequences (i.e., items with only positive feedback), building on the observations from Section 4, We preserve the complete interaction sequence for each user, explicitly retaining all historical items along with their associated preference feedback. We use explicit ratings or ratings converted from implicit feedback to construct a hierarchical preference profile. Following prior work Chen et al. (2024), the input prompts are composed of the following parts:
@@ 5.1 Comprehensive and Structured Preference Feedback as the Foundation
We prepend the history-specific prefixes (e.g., "Given the user’s recent viewing and rating history") and candidate-specific prefixes (e.g., "recommend a movie they’ll likely watch next and rate generously from following candidates") to their respective sequences.
The three prompt components are concatenated as the final textual input $\mathbf{x}_{u}$ to the LMs. Concrete examples are demonstrated in Appendix C.
5.2 Adaptive Reward Margin Emulates Complex Human Preferences
Current DPO-based methods simplify preference modeling to maximizing the reward difference between pairwise preferred and dispreferred responses/items, exposing them to two key limitations: (i) Neglecting preference hierarchy, where in reality, users may strongly prefer certain items while only slightly prefer others, compared to either the same or different negative items; (ii) Neglecting contextual factors, among which we specifically consider that users typically prioritize immediate satisfaction over delayed rewards.
@@ 5.2 Adaptive Reward Margin Emulates Complex Human Preferences
where $\mathbf{y}_{p}$ is preferred over $\mathbf{y}_{d}$, $s_{p}$ and $s_{d}$ are their structured preference score respectively, $\lambda$ controls the margin’s magnitude, and $\Delta_{t_{p}}=t_{p}^{+}-t$ indicates the time latency of the interaction. In this work, we set $\phi\left(s,\Delta_{t}\right)=s/\left(\Delta_{t}\right)^{0.5}$. Note that the choice of score function is customizable as long as it reflects the above preference rules. That is, $\phi\left(s,\Delta_{t}\right)\propto s/\left(\Delta_{t}\right)^{\alpha}$, where $\alpha>0$ indicates the temporal decay factor. For dispreferred items from either negative sampling or historical interactions where no user-assigned feedback is available, we set a default preference score and time latency to facilitate the training. More details about the default value can be found in Section 6.
5.3 Complex Preference Pattern Emerges from Adaptive Reward Margins
@@
where $r(\cdot)$ is the reward function. We pair each preferred item with multiple dispreferred items, and leverage the Plackett-Luce (PL) model Plackett (1975); Luce (1959) to generalize pairwise comparisons to a list-wise ranking framework. Formally, given the prompt $x_{u}^{t}$ encompassing all the historical interactions of user $u$, a candidate set $\mathcal{C}$ containing $K$ items (one preferred item and $K-1$ dispreferred items), and a permutation $\sigma$ representing the predicted ranking of these candidates based on user preference for the next item (denote $\sigma(j)$ as the item ranked at position $j$), the probability of observing the candidates’ preference ranked as $[\mathbf{y}_{\sigma(1)},\mathbf{y}_{\sigma(2)},\ldots,\mathbf{y}_{\sigma(K)}]$ is:
@@
Model Type Bkbn Method MovieLens Amazon-Books BeerAdvocate Steam LastFM HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio Feedback Type Explicit Feedback Implicit Feedback Trad. - GRU4Rec 0.2664 1.0000 0.1310 1.0000 0.3708 1.0000 0.4584 1.0000 0.6630 1.$0$ Caser 0.2714 1.0000 0.1538 1.0000 0.3757 1.0000 0.4394 1.0000 0.6716 1.$0$ SASRec 0.2671 1.0000 0.1559 1.0000 0.3800 1.0000 0.4587 1.0000 0.6659 1.0000 LLM LLaMA3-8B LLaMA3 0.0929 0.7351 0.0654 0.6165 0.0686 0.6617 0.0852 0.8672 0.1264 0.6147 SFT 0.2478 0.9985 0.4447 0.9974 0.2645 0.9936 0.3122 0.9990 0.5076 1.0000 DPO 0.2809 0.9970 0.5049 0.9887 0.4412 0.9875 0.3340 0.9980 0.5719 1.0000 SimPO 0.2974 0.9725 0.5129 0.9564 0.4020 0.9250 0.3401 0.9766 0.5759 0.9419 S-DPO 0.2902 0.9983 0.5065 0.9880 0.4698 0.9903 0.3588 0.9990 0.5719 0.9990 RecPO 0.3451 0.9969 0.5802 0.9851 0.5771 0.9887 0.4672 0.9985 0.6830 0.9959 Qwen-7B Qwen 0.1204 0.7471 0.1013 0.7194 0.0583 0.4223 0.1477 0.6293 0.2148 0.6860 SFT 0.2060 0.9983 0.3659 0.9967 0.2044 0.9849 0.2081 0.9950 0.3119 0.9969 DPO 0.2610 0.9983 0.4412 0.9930 0.2600 0.9724 0.2457 0.9960 0.4046 0.9969 SimPO 0.2888 0.9531 0.4644 0.9880 0.4044 0.9529 0.3706 0.9940 0.5209 0.9796 S-DPO 0.2706 0.9957 0.4623 0.9910 0.3253 0.9798 0.3062 0.9970 0.4495 0.9959 RecPO 0.3446 0.9896 0.5307 0.9880 0.4320 0.9729 0.4143 0.9912 0.5973 0.9980

### v2

# What Makes LLMs Effective Sequential Recommenders? A Study on Preference Intensity and Temporal Context

@@ Abstract
Sequential recommendation systems aspire to profile users by interpreting their interaction histories, echoing how humans make decisions by weighing experience, relative preference strength, and situational relevance. Yet, existing large language model (LLM)-based recommenders often fall short of mimicking the flexible, context-aware decision strategies humans exhibit, neglecting the structured, dynamic, and context-aware mechanisms fundamental to human behaviors. To bridge this gap, we propose RecPO, a preference optimization framework that models structured feedback and contextual delay to emulate human-like prioritization in sequential recommendation. RecPO exploits adaptive reward margins based on inferred preference hierarchies and temporal signals, enabling the model to favor immediately relevant items and to distinguish between varying degrees of preference and aversion. Extensive experiments across five real-world datasets demonstrate that RecPO not only yields performance gains over state-of-the-art baselines, but also mirrors key characteristics of human decision-making: favoring timely satisfaction, maintaining coherent preferences, and exercising discernment under shifting contexts. Code: https://anonymous.4open.science/r/RecPO-020A/
What Makes LLMs Effective Sequential Recommenders?
A Study on Preference Intensity and Temporal Context
@@ 1 Introduction
Large language models (LLMs) are increasingly being adapted for sequential recommendation Harte et al. (2023); Li et al. (2023); Yang et al. (2024); Bao et al. (2023); Zhang et al. (2023), where the task is to predict the next item a user will interact with based on their historical behaviors. Unlike traditional recommenders Hidasi (2016); Kang and McAuley (2018); Tang and Wang (2018), LLM-based systems leverage semantic understanding and reasoning capabilities to model user preferences from textual interaction histories.
Figure 1: Human behaviors involve trade-offs among preference intensity, satisfaction delay, effort, and risk—factors largely overlooked in current LLM-based preference modeling.
Current approaches predominantly rely on preference alignment techniques such as DPO Rafailov et al. (2024) and its variants Chen et al. (2024); Meng et al. (2024); Amini et al. (2024), which treat all preferences uniformly through binary pairwise comparisons. This binary abstraction, while effective for general language tasks, misaligns with human decision-making: humans exhibit graded preferences (strongly love vs. mildly like) and temporal sensitivity (immediate vs. delayed satisfaction), as illustrated in Figure 1. Such graded and temporally-aware patterns are pervasive in human behavior Astington and Jenkins (1995), yet remain largely unmodeled in current LLM-based recommenders. This raises a critical question: what specific factors in preference data enable LLMs to capture these nuanced human behaviors for recommendation?
We investigate this question through systematic empirical study: Our proof-of-concept experiment (§ 3) reveals that incorporating comprehensive feedback (including negative interactions) and structured preference signals (e.g., ratings) substantially improves performance, suggesting binary modeling discards critical information. Through controlled ablations, we identify two key factors: (1) preference intensity—the graded strength of user affinity or aversion, and (2) temporal context—the immediacy of satisfaction. These factors, though well-established in behavioral economics and cognitive science, have been largely overlooked in LLM preference alignment for recommendation.
Building on these insights, we introduce RecPO, a preference optimization framework that operationalizes both preference intensity and temporal context through adaptive reward margins. Unlike prior work that only uses uniform margins across all preference pairs, RecPO discovers more fine-grained preference signal to model human preference: (i) the graded preference scores of items (e.g., 5-star vs. 3-star ratings), and (ii) their temporal distance from the current decision point. This enables the LLM to mirror key characteristics exhibit in human preference behaviors. Our contributions are threefold:
  • We systematically demonstrate that preference intensity and temporal context are fine-grained factors for LLM-based preference modeling in recommendation, challenging the prevailing binary preference paradigm (§ 3).
  • By incorporating these factors through adaptive reward margins derived from preference intensity and temporal context into LLM, we provide RecPO for sequential recommendation (§ 4).
  • Through experiments on five datasets with both explicit and implicit feedback, we show that RecPO improves accuracy with behavioral characteristics aligned with human preference: prioritizing timely satisfaction and maintaining preference coherence under shifting contexts (§ 5).
@@ 2 Preliminaries
where $\pi_{\text{ref}}$ is commonly set to $\pi_{\text{SFT}}$, $\pi_{\theta}$ is the aligned model, and $\beta$ is a hyperparameter. Building upon DPO, a recent effort named S-DPO Chen et al. (2024) has been proposed specifically for LLM-based recommenders. They pair each positive item with multiple negative items generated by random sampling as preference data, and revise the alignment objective as:
@@
Figure 3: Illustrations for preference learning frameworks with binary and enriched preference: the prior assumes binary distinction in preference, while the latter enriches preference distinction with preference intensity and temporal context ($\delta$ indicates the enrichment).
3 What Do Current Methods Overlook? A Proof-of-Concept Investigation
Current LLM-based recommenders, including S-DPO Chen et al. (2024), typically filter out negative feedback items from user histories and discard structured preference signals (e.g., ratings), treating all remaining items uniformly. But does this practice discard critical information?
To investigate, we design a proof-of-concept experiment that varies two dimensions of user feedback: comprehensiveness (whether negative interactions are retained) and structure (whether graded preference signals are provided). We devise four-tier input configurations that progressively integrate preference signals: (i) Filtered Items: Excluding negative feedback items and no explicit ratings are provided, mimicking S-DPO’s setup; (ii) Full Items: Retaining all historical items, yet no explicit ratings are provided; (ii) Filtered Items + Rating: Providing explicit ratings yet excluding negative-feedback items; (iv) Full Items + Rating: Retaining all items and their corresponding explicit ratings.
We fine-tune LLaMA3-8B on MovieLens and Amazon-Books ( described § 5.1) using the four input configurations. The experimental results are reported using Hit Ratio@1 (see § 5.1, where higher values indicate better performance) and are shown in Figure 2.
Key Findings.
Three striking patterns emerge: (1) Comprehensive feedback matters: Comparing Filtered Items + Rating vs. Full Items + Rating, retaining negative interactions consistently improves performance. While counterintuitive—why include disliked items when predicting the next favorable item?—this suggests that aversion modeling is crucial for accurate preference profiling. (2) Structured signals matter: Comparing Full Items vs. Full Items + Rating, adding structured preference signals (ratings) substantially boosts performance. Without ratings, Full Items even underperforms Filtered Items, as negative items become noise without explicit annotations. (3) Both factors are complementary: The best performance comes from combining both—comprehensive feedback with structured signals.
These results reveal that current methods overlook two critical aspects: the graded strength of preferences (intensity) and the distinction between positive and negative feedback. However, what specific factors enable LLMs to leverage this richer information? This motivates our investigation into preference intensity and temporal context as the underlying mechanisms.
@@ 4 Methodology
We first lay out the prompt design that establishes the foundation for preference modeling in LLM-based recommendations. We then introduce RecPO, a novel preference optimization framework for sequential recommendation that dynamically calibrates reward margins between pairwise preference data based on contextualized and structured preference feedback, shown in Figure 3.
4.1 Operationalizing Comprehensive and Structured Feedback
Unlike existing approaches Liao et al. (2024); Chen et al. (2024) which remove items with negative feedback from interaction histories to construct homogeneous sequences (i.e., items with only positive feedback), building on the observations from § 3, We preserve the complete interaction sequence for each user, explicitly retaining all historical items along with their associated preference feedback. We use explicit ratings or ratings converted from implicit feedback to construct a hierarchical preference profile. Following prior work Chen et al. (2024), the input prompts are composed of the following parts:
@@ 4.1 Operationalizing Comprehensive and Structured Feedback
We prepend the history-specific prefixes (e.g., "Given the user’s recent viewing and rating history") and candidate-specific prefixes (e.g., "recommend a movie they’ll likely watch next and rate generously from following candidates") to their respective sequences. The three prompt components are concatenated as the final textual input $\mathbf{x}_{u}$ to the LMs. Concrete examples are demonstrated in Appendix C.
4.2 Modeling Preference Intensity and Temporal Context
Current preference learning methods simplify preference modeling to maximizing the reward difference between pairwise preferred and dispreferred responses/items, exposing them to two key limitations: (i) Neglecting preference intensity, where in reality, users may strongly prefer certain items while only slightly prefer others, compared to either the same or different negative items; (ii) Neglecting temporal context,where users typically prioritize immediate satisfaction over delayed rewards.
@@ 4.2 Modeling Preference Intensity and Temporal Context
where $\mathbf{y}_{p}$ is preferred over $\mathbf{y}_{d}$, $s_{p}$ and $s_{d}$ are their structured preference score respectively, $\lambda$ controls the margin’s magnitude, and $\Delta_{t_{p}}=t_{p}^{+}-t$ indicates the time latency of the interaction. In this work, we set $\phi\left(s,\Delta_{t}\right)=s/\left(\Delta_{t}\right)^{0.5}$. Note that the choice of score function is customizable as long as it reflects the above preference rules. That is, $\phi\left(s,\Delta_{t}\right)\propto s/\left(\Delta_{t}\right)^{\alpha}$, where $\alpha>0$ indicates the temporal decay factor. For dispreferred items from either negative sampling or historical interactions where no user-assigned feedback is available, we set a default preference score and time latency to facilitate the training. More details about the default value can be found in § 5.
4.3 Deriving the Preference Alignment Objective
@@
where $r(\cdot)$ is the reward function. We pair each preferred item with multiple dispreferred items, and leverage the Plackett-Luce (PL) model Plackett (1975); Luce (1959) to generalize pairwise comparisons to a list-wise ranking framework. Formally, given the prompt $x_{u}^{t}$ encompassing all the historical interactions of user $u$, a candidate set $\mathcal{C}$ containing $K$ items (one preferred item and K1K-1 dispreferred items), and a permutation $\sigma$ representing the predicted ranking of these candidates based on user preference for the next item (denote $\sigma(j)$ as the item ranked at position $j$), the probability of observing the candidates’ preference ranked as $\mathbf{y}_{\sigma(1)},\mathbf{y}_{\sigma(2)},\ldots,\mathbf{y}_{\sigma(K)}]$ is:
@@
Model Type Bkbn Method MovieLens Amazon-Books BeerAdvocate Steam LastFM HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio HR@1 ValidRatio Feedback Type Explicit Feedback Implicit Feedback Trad. - GRU4Rec 0.2664 1.0000 0.1310 1.0000 0.3708 1.0000 0.4584 1.0000 0.6630 1.0000 - Caser 0.2714 1.0000 0.1538 1.0000 0.3757 1.0000 0.4394 1.0000 0.6716 1.0000 - SASRec 0.2671 1.0000 0.1559 1.0000 0.3800 1.0000 0.4587 1.0000 0.6659 1.0000 LLM LLaMA3-8B LLaMA3 0.0929 0.7351 0.0654 0.6165 0.0686 0.6617 0.0852 0.8672 0.1264 0.6147 SFT 0.2478 0.9985 0.4447 0.9974 0.2645 0.9936 0.3122 0.9990 0.5076 1.0000 DPO 0.2809 0.9970 0.5049 0.9887 0.4412 0.9875 0.3340 0.9980 0.5719 1.0000 SimPO 0.2974 0.9725 0.5129 0.9564 0.4020 0.9250 0.3401 0.9766 0.5759 0.9419 S-DPO 0.2902 0.9983 0.5065 0.9880 0.4698 0.9903 0.3588 0.9990 0.5719 0.9990 RecPO 0.3451 0.9969 0.5802 0.9851 0.5771 0.9887 0.4672 0.9985 0.6830 0.9959 Qwen-7B Qwen 0.1204 0.7471 0.1013 0.7194 0.0583 0.4223 0.1477 0.6293 0.2148 0.6860 SFT 0.2060 0.9983 0.3659 0.9967 0.2044 0.9849 0.2081 0.9950 0.3119 0.9969 DPO 0.2610 0.9983 0.4412 0.9930 0.2600 0.9724 0.2457 0.9960 0.4046 0.9969 SimPO 0.2888 0.9531 0.4644 0.9880 0.4044 0.9529 0.3706 0.9940 0.5209 0.9796 S-DPO 0.2706 0.9957 0.4623 0.9910 0.3253 0.9798 0.3062 0.9970 0.4495 0.9959 RecPO 0.3446 0.9896 0.5307 0.9880 0.4320 0.9729 0.4143 0.9912 0.5973 0.9980